OpenStudy (anonymous):

the populationgrowth model is y=ce^rt where Y is FV and C is PV . ind the rate of growth in %ages if the population growth in 2001 was 2.3 mill and 2010 3.1 mill

OpenStudy (anonymous):

do you know how to start?

OpenStudy (anonymous):

well 3.1m = 2.3 (? ^ r(9) ?

OpenStudy (anonymous):

okay we have y=ce^(rt) we don't know c or r. we are trying to find r. so 2.3=ce^(2001r) [1] and 3.1=ce^(2010r) [2] to get rid of c we are going to divide [1] by [2] 2.3/3.1=e^(2001r)/e^(2010r) 0.7419=e^(2001r-2010r) 0.7419=e^(-9r) ln(0.7419)=ln(e^(-9r)) -0.2985=-9r r=-0.2985/-9 =0.0332

OpenStudy (anonymous):

thankss